OpenStudy (anonymous):

solve x-y=6. please? step by step :>

OpenStudy (anonymous):

y=6-x

OpenStudy (anonymous):

plug that into the previous equation

OpenStudy (anonymous):

so x-(6-x)=6

OpenStudy (anonymous):

x-6+x=6

OpenStudy (anonymous):

x+x=12

OpenStudy (anonymous):

2x=12

OpenStudy (anonymous):

x=6, i'm almost POSITIVE that's the answer

OpenStudy (anonymous):

and so y would equal 0?

OpenStudy (anonymous):

What is question and what u answered